4.5 g of copper reacts with 1.50 L of 0.225 M nitric acid according to the following reaction. If the percent yield is 89.0%, what is the actual volume of nitrogen dioxide gas produced at 84.0 kPa and 42.0 degrees Celsius?
1
Expert's answer
2020-12-17T07:38:40-0500

4 HNO3(l) + Cu(s) ==> Cu(NO3)2(aq) + 2 NO2(g) + 2 H2O(l)

Moles of Cu=4.5/63.546=0.0708moles

Moles of HNO3=1.5×0.225

                        =0.3375 moles

The moles that were all consumed in reaction are 0.0708moles of Cu

Moles of NO2= (0.0708)4

=0.1416molesV2=P1V1T2/T1P2

At room temperature the 1mole=24dm3

                                              =0.1416 dm3

                                                 =24×0.1416

                                                   =3.3984 dm3

T2=315K

1atm=101325Pa

%Yield of NO2=0.89 ×3.3984

=3.02563 dm3

V2=315×101325×3.0246/ (298×8400)

V2=38.57 dm3